OpenStudy (anonymous):

Please Help!! Part 1: Is the following equation true? (2 points) Part 2: Use complete sentences to explain the properties used in making your decision. Equation is below.

OpenStudy (anonymous):

\[3\log _{2}x + 1/2 \log _{2}y-3 \log _{2}z=\log _{2}x+3\sqrt{y}/z ^{3}\]

jhonyy9 (jhonyy9):

step by step please

jhonyy9 (jhonyy9):

how you have got this result ?

OpenStudy (anonymous):

no i'm confused on how to do it.

OpenStudy (jit4won):

It's simple

OpenStudy (anonymous):

please explain then how to do it.

OpenStudy (jit4won):

Use properties carefully